OpenStudy (trojanpoem):

if you knew that: Pw = VI the dimensions are: a) ML^2A^-1T^-3 b) MLA^-1T^-2 c)ML^-2AT^-3 D)ML^2A^-1T^-2

OpenStudy (trojanpoem):

I tried V = W/Q pw = WI/Q Q = it pw = w/t Dimensions: ML^2T^-3 What's wrong ?

OpenStudy (michele_laino):

we have: volt* Amperes= (joule/ coulomb)*amperes=\[ = \frac{{joules}}{{amperes \cdot \sec }} \cdot amperes = \frac{{joules}}{{\sec }}\]

OpenStudy (michele_laino):

now, using dimensions, we can write this: \[\frac{{work}}{{time}} = \frac{{ML{T^{ - 2}}L}}{T} = M{L^2}{T^{ - 3}}\]

OpenStudy (trojanpoem):

That's none of the choices :/

OpenStudy (michele_laino):

what is A?

OpenStudy (trojanpoem):

A is the dimension of Electric current

OpenStudy (trojanpoem):

Is the question a typo ?

OpenStudy (michele_laino):

I think so, since as we can see from my steps above, amperes cancel out, since we have on amperes at numerator , and another one at denominator, so the right expression can be this one: \[\frac{{work}}{{time}} = \frac{{ML{T^{ - 2}}L}}{T} = M{L^2}{T^{ - 3}}{A^0}\]

OpenStudy (michele_laino):

one*

OpenStudy (trojanpoem):

Yeah. I believe so but was making sure :) thanks again

OpenStudy (michele_laino):

:)